OpenStudy (anonymous):

If a uranium atom has a mass number of 234, how many neutrons are in its nucleus? Not math I know... >:P But I'd REALLY appreciate the help!!! :)

OpenStudy (anonymous):

\begin{array}l\color{red}{\normalsize\text{C}}\color{orange}{\normalsize\text{H}}\color{#9c9a2e}{\normalsize\text{E}}\color{green}{\normalsize\text{M}}\color{blue}{\normalsize\text{I}}\color{purple}{\normalsize\text{S}}\color{purple}{\normalsize\text{T}}\color{red}{\normalsize\text{R}}\color{orange}{\normalsize\text{Y}}\color{#9c9a2e}{\normalsize\text{ }}\color{green}{\normalsize\text{a}}\color{blue}{\normalsize\text{n}}\color{purple}{\normalsize\text{d}}\color{purple}{\normalsize\text{ }}\color{red}{\normalsize\text{P}}\color{orange}{\normalsize\text{H}}\color{#9c9a2e}{\normalsize\text{Y}}\color{green}{\normalsize\text{S}}\color{blue}{\normalsize\text{I}}\color{purple}{\normalsize\text{C}}\color{purple}{\normalsize\text{S}}\color{red}{\normalsize\text{}}\end{array}

OpenStudy (anonymous):

Yea....When there are as many ACTIVE people in the chem or physics room THEN I'll post it there.

OpenStudy (anonymous):

NO FAIR!! I cant type all cool like that >:P lol

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

Teach meeeeeeeeeeeeeeeeeeeeee

OpenStudy (anonymous):

Teach you? Didn't amistre64 give you the program to do this?

OpenStudy (anonymous):

nope >:P

OpenStudy (anonymous):

U234 Atomic Number -92 Mass number-234 No: of neutrons=234-92=142